# ALMA: a mathematics-driven approach for determining tuning parameters in generalized LASSO problems, with applications to MRI

### v1

@@ 5. Solve
Note that the scalar product $a^{T}b$, $a,b\in\mathbb{R}^{n}$, must be replaced with $Re(a^{T}b)$ when $a,b\in\mathbb{C}^{n}$, where $Re$ denotes the real part. Moreover, for non-Cartesian MRI trajectories, the scalar product that shall be considered is $\langle a,b\rangle=Re(a^{T}Hb)$, with $H$ Hermitian positive-definite matrix encoding the non-Cartesian gridding of the k-space.
In the previous paragraphs we observed that finding an ALM is a matter of tracing the tangent line in 0 to the epigraph $\mathcal{A}=\{(u,t)\in\mathbb{R}^{2}:u\geq\|Ax-b\|_{2}^{2}/2-\eta^{2}/2,\ t\geq TV(x)/2\ for\ some\ x\in\mathbb{R}^{n}\}$, where $\eta=\|Af-b\|_{2}$, being $f\in\mathbb{R}^{384\times 384}$ the $384\times 384$ Shepp-Logan phantom. We pointed out that outlining $\mathcal{A}$ is technically difficult due to the unfeasible dimensionality: in principle, for every $x\in\mathbb{R}^{n}$ ($n=384^{2}$), once the point $(u,t)=(u(x),t(x))=1/2\cdot(\|Ax-b\|_{2}^{2}-\eta^{2},TV(x))$ is computed, one has that all the points of the first quadrant centered in $(u(x),t(x))$ belong to $\mathcal{A}$. Clearly, it would be enough to compute $(u(x),t(x))$ only for $x\in\mathbb{R}^{n}$ such that $(u(x),t(x))$ belongs to the boundary $\partial\mathcal{A}$ of $\mathcal{A}$, but we do not have access to those points. On top of that, computing $(u(x),t(x))$ for a fixed $x$ is computationally expensive, because of the measurement operator $A$. However, a meaningful family of points $(u(x),t(x))$ that would be enough to approximate the tangent line in 0 to its boundary can be found as follows:
@@
    Choose $x\in\mathbb{R}^{n}$ so that the corresponding $(u(x),t(x))$ is as far to the left of $\mathcal{A}$ as possible. Clearly, this task is accomplished by any minimizer of $\phi(x)=\|Ax-b\|_{2}^{2}$, which consists of iterative SENSE reconstructions. Let us call this point $x^{\#}$.
@@
    Choose $\lambda^{(1)}=-1/m^{(1)}$, where $m^{(1)}$ is the slope of the tangent and reconstruct an image using $\lambda^{(1)}$ as tuning parameter for TV-LASSO. Call this image $x^{(1)}$.
@@
III-C Image quality metrics
@@
  • The mSSIM is an extension of the structural similarity index, designed to assess the quality of reconstructions across various scales in a manner that approximates human perception, c.f. [14, 15]. It compares the brightness, contrast, and structural details of reconstructions with ground truth images, assigning values on a scale from 0 to 1, where a score of 1 indicates optimal similarity. Good quality reconstructions typically correspond to mSSIM values of $\geq 0.9$. In the present work, the mSSIM is computed via the command multissim(I,Iref), where Iref is the reference image (the Shepp-Logan brain phantom) and I is the image to be assessed.
  • The pSNR quantifies noise corruption of compressed images, independently on the quality as perceived by human vision and good visual quality requires pSNR to be at least 30dB, c.f. [16].
  • CJV measures the presence of intensity non-uniformity (INU) artifacts in MRI, c.f. [17, 18]. In the current paper, we use it as a measure of artifact bias in reconstructions. Lower values of CJV indicate better quality of MR images in terms of artifacts. CJV is defined based on the intensity difference between grey and white matter. Fig. 5 illustrates the masks corresponding to grey matter (B) and white matter (C) in the Shepp-Logan phantom, used for the computation of the CJV, as inspired by [19].
A Refer to caption   B Refer to caption   C Refer to caption
Figure 5: The Shepp-Logan brain phantom (A). Grey matter (B) and white matter (C) masks for CJV computation.
@@ A Refer to caption   B Refer to caption   C Refer to caption
III-D Data analysis
The mSSIM, pSNR, and CJV, along with their respective relative versions, serve as quantitative measures for assessing the quality of reconstructions. Additionally, the (unique, in our experiments) tuning parameters $\lambda_{mSSIM},\lambda_{pSNR}$ and $\lambda_{CJV}$ which optimize each metric can be considered. Specifically, $\lambda_{mSSIM}=\arg\max\{mSSIM(\lambda):\lambda\geq 0\}$, $\lambda_{pSNR}=\arg\max\{pSNR(\lambda):\lambda\geq 0\}$ and $\lambda_{CJV}=\arg\min\{CJV(\lambda):\lambda\geq 0\}$. The ratios $\lambda_{mSSIM}/\lambda_{ALM}$, $\lambda_{pSNR}/\lambda_{ALM}$ and $\lambda_{CJV}/\lambda_{ALM}$ express the distance between optimal tuning parameters (with respect to the metrics) and ALMs in terms of the orders of magnitude of the corresponding ratios. We term the ratio $\lambda_{mSSIM}/\lambda$ the magnitude ratio corresponding to $\lambda$ (with respect to mSSIM), and similar notations are reserved for the other metrics. For every pair $({UR}_{\%},{NL}_{\%})$ and every reconstruction, we calculate the mSSIM, pSNR and CJV, alongside their corresponding magnitude rations. Within each fixed pair $({UR}_{\%},{NL}_{\%})$, we employ violin plots to illustrate the distributions of mSSIM, pSNR and CJV across the 50 runs, for analysis. Additionally, shaded error bars for the functions $mSSIM(\lambda/\lambda_{ALM})$, $pSNR(\lambda/\lambda_{ALM})$ and $CJV(\lambda/\lambda_{ALM})$ are incorporated to evaluate the optimality of ALMs as tuning parameters. A value of $\lambda_{mSSIM}/\lambda_{ALM}=1$ signifies that $\lambda_{ALM}$ is the optimal tuning parameter with respect to mSSIM. Similar interpretations apply for the other magnitude ratios. For a comparison, we compute the L-curve tuning parameter $\lambda_{L}$, and execute reconstructions using TV-LASSO with this parameter. Subsequently, we compute $mSSIM(\lambda_{L})$, $pSNR(\lambda_{L})$ and $CJV(\lambda_{L})$, along with the corresponding magnitude ratios pertaining to $\lambda_{L}$ with respect to these three metrics.
III-E The dataset
@@
IV Results
@@
IV-A Analysis of convergence
A Refer to caption   B Refer to caption
Figure 7: Convergence analysis. Histograms of the number of iterations needed for ALMA’s convergence across undersampling rates (A) and noise levels (B). The number of bins for each histogram is the ceiling of the square root of the number of points, i.e., $\lceil\sqrt{150}\rceil=13$. Observe that the number of iterations for ALMA to stop is mostly concentrated in the interval $[0,10]$ independently of the noise level and the undersampling rate, while the standard deviations increase with the undersampling rate and the noise percentage. This demonstrates that the number of iterations required for ALMA to converge is relatively low even when processing larger amounts of information ($20\%$ undersampling) and higher noise percentages ($7\%$ noise).
For fixed noise level and undersampling rate, ALMA approximates an ALM by conducting reconstructions, updating the ALM approximation at each iteration. We have selected two stopping criteria:
@@ A Refer to caption   B Refer to caption
Note that the second criterion serves as a convergence criterion for ALMA. This means that if it is satisfied, ALMA not only stops but actually converges (i.e., the sequence $(\lambda^{(k)})_{k}$ of ALMs at step $k$ converges). Observe that in all of our experiments, the sequence of the ALMs is eventually constant. Therefore, in what follows, the term convergence will refer to ALMA stopping due to the fulfillment of the second stopping criterion.
@@
Through empirical analysis, we investigated the convergence of ALMA across various noise levels and undersampling rates. As aforementioned, ALMA converges in a finite time for every undersampling rate and noise level.
@@
The histograms in fig. 7 (A) illustrates the number of iterations required for the convergence of ALMA across different noise levels and undersampling rates. Both histograms follow Gaussian distributions that become less concentrated around 0 as the undersampling rate increases (fig. 7 (A)) or as the noise level rises (fig. 7 (B)). In both scenarios, the increase in the number of iterations can be attributed to the amount of information processed by ALMA, which grows with the number of sampling points and the noise level. Let us delve deeper into these cases. On average, ALMA stops after $7.2089\pm 2.9773$ iterations (see fig. 9). For the $10\%$ sampling rate, both the average number of iterations and the standard deviation increase with the noise level. A similar trend is observed for the $15\%$ undersampling rate, as detailed in fig. 9. However, the $20\%$ sampling rate exhibits a distinct pattern: the average number of iterations increases from $3\%$ to $5\%$ noise levels and reaches its minimum at a noise level of $7\%$.
IV-B Performance of ALMA with respect to mSSIM
ARefer to captionBRefer to captionCRefer to caption
Figure 8: Shaded error bar of the mSSIM (A), pSNR (B) and CJV (C) as functions of $\lambda/\lambda_{ALM}$, the shades representing the corresponding standard deviations. The values 1 on the horizontal axes correspond to $\lambda=\lambda_{\text{ALM}}$.
@@ IV-B Performance of ALMA with respect to mSSIM
Images reconstructed utilizing the ALMs computed by ALMA consistently exhibit an average $mSSIM\geq 0.99$, irrespective of the noise level or the undersampling rate (see fig. 9). The violin plots for the mSSIM are displayed in figure 6 (color yellow).
@@
Across noise levels and undersampling rates, the tuning parameter that optimizes the mSSIM ($\lambda_{mSSIM}$) is around half of the ALM ($\lambda_{ALM}$), i.e. $\lambda_{mSSIM}\approx 0.52\cdot\lambda_{ALM}$ (see fig. 9). However, even if $\lambda_{ALM}$ does not maximize the mSSIM, the points $(1,mSSIM(1))$, which correspond to the mSSIM of the reconstructions obtained with ALMs, always lie on the plateau of the graph of the function $mSSIM(\lambda/\lambda_{ALM})$, indicating that $\lambda_{ALM}$ is in the range of tuning parameters corresponding to almost-optimal mSSIM values (see fig. 8 (A)).
IV-C Performance of ALMA with respect to pSNR
@@
Reconstructions obtained using ALMA exhibit high pSNR for every noise level and undersampling rate. As expected the pSNR decreases across noise levels. On average, the pSNR is $\geq 40$dB for reconstructions of images corresponding to $15\%$ and $20\%$ undersampling rates, whereas the average pSNR of reconstructions with $10\%$ undersampling rate is at least 35dB (see fig. 9). Other than demonstrating the quality of reconstructions obtained via ALMA in terms of pSNR, this reflects the fact that solving TV-LASSO with ALMs as tuning parameters tend to produce highly regularized images. Again, across noise levels and undersampling rates, the tuning parameter that optimizes the pSNR ($\lambda_{pSNR}$) tends to be approximately 0.47 times the corresponding ALM, i.e. $\lambda_{pSNR}\approx 0.47\cdot\lambda_{ALM}$ (see fig. 9). Compared to the graphs $mSSIM(\lambda/\lambda_{ALM})$, the plateau of the functions $pSNR(\lambda/\lambda_{pSNR})$ near their maxima is less pronounced, and the points $(1,pSNR(1))$, corresponding to the pSNR of reconstructions obtained with ALMs, are shifted to the right of their peaks (see fig. 8 (B)). In conclusion, ALMs perform almost-optimally with respect to pSNR.
IV-D Performance of ALMA with respect to CJV
@@
Except for the worst-case scenario $UR_{\%}=10\%$, $NL_{\%}=7\%$, the reconstructions obtained by ALMA display CJV values no larger than 0.05, showing that ALMA performs well with respect to CJV as well (see fig. 9). The tuning parameter that minimizes the CJV ($\lambda_{CJV}$) is on average 0.45 times the corresponding ALM, i.e. $\lambda_{CJV}\approx 0.45\cdot\lambda_{ALM}$ (see fig. 9). However, ALMs still perform almost optimally with respect to CJV, and the points $(1,CJV(1))$, corresponding to the CJV of reconstructions obtained with ALMs, are shifted to the right of their minima, in accordance with the behavior of the graphs $pSNR(\lambda/\lambda_{CJV})$ (see fig. 8 (C)).
IV-E Comparison with the L-curve parameter

### v2

@@ 5. Solve
Note that the scalar product $a^{T}b$, $a,b\in\mathbb{R}^{n}$, must be replaced with $Re(a^{T}b)$ when $a,b\in\mathbb{C}^{n}$, where $Re$ denotes the real part [12]. Moreover, for non-Cartesian MRI trajectories, the scalar product that shall be considered is $\langle a,b\rangle=Re(a^{T}Hb)$, with $H$ Hermitian positive-definite matrix encoding the non-constant density of the k-space sampling.
In the previous paragraphs we observed that finding an ALM is a matter of tracing the tangent line in 0 to the epigraph
[…]
$u\geq\|Ax-b\|_{2}^{2}/2-\eta^{2}/2,\ t\geq TV(x)/2,$
$\ \text{for \ some}\ x\in\mathbb{C}^{n}\}$
where $\eta=\|Af-b\|_{2}$, being $f\in\mathbb{R}^{384\times 384}$ the $384\times 384$ Shepp-Logan phantom. We pointed out that outlining $\mathcal{A}$ is technically difficult due to the unfeasible dimensionality: in principle, for every $x\in\mathbb{C}^{n}$ ($n=384^{2}$), once the point
$(u,t)=(u(x),t(x))=1/2\cdot(\|Ax-b\|_{2}^{2}-\eta^{2},TV(x))$
is computed, one has that all the points of the first quadrant centered in $(u(x),t(x))$ belong to $\mathcal{A}$. Clearly, it would be enough to compute $(u(x),t(x))$ only for $x\in\mathbb{C}^{n}$ such that $(u(x),t(x))$ belongs to the boundary $\partial\mathcal{A}$ of $\mathcal{A}$, but we do not have access to those points. On top of that, computing $(u(x),t(x))$ for a fixed $x$ is computationally expensive, because of the measurement operator $A$. However, a meaningful family of points $(u(x),t(x))$ that would be enough to approximate the tangent line in 0 to its boundary can be found as follows:
@@
    Choose $x\in\mathbb{C}^{n}$ so that the corresponding $(u(x),t(x))$ is as far to the left of $\mathcal{A}$ as possible. Clearly, this task is accomplished by any minimizer of $\phi(x)=\frac{1}{2}\|Ax-b\|_{2}^{2}$. Let us call this point $x^{\#}$.
@@
    Choose $\lambda^{(1)}=-1/m^{(1)}$, where $m^{(1)}$ is the slope of the tangent and reconstruct an image using $\lambda^{(1)}$ as tuning parameter for TV-LASSO. Call this image $x^{(1)}$. Note that this step consist in performing a reconstruction. This is why ALMA can be considered as an iterative reconstruciton while adapting $\lambda$ until convergence.
@@
2.2.3 Image quality metrics
@@
  • The mSSIM is an extension of the structural similarity index, designed to assess the quality of reconstructions across various scales in a manner that approximates human perception,[18, 19]. It compares the brightness, contrast, and structural details of reconstructions with ground truth images, assigning values on a scale from 0 to 1, where a score of 1 indicates optimal similarity. Good quality reconstructions typically correspond to mSSIM values of $\geq 0.9$. In the present work, the mSSIM is computed via the command multissim(I,Iref), where Iref is the reference image (the Shepp-Logan brain phantom) and I is the image to be assessed.
  • The pSNR quantifies noise corruption of compressed images, independently on the quality as perceived by human vision and good visual quality requires pSNR to be at least 30dB, [20].
  • CJV measures the presence of intensity non-uniformity (INU) artifacts in MRI [21, 22]. In the current paper, we use it as a measure of artifact bias in reconstructions. Lower values of CJV indicate better quality of MR images in terms of artifacts. CJV is defined based on the intensity difference between grey and white matter, as inspired by Jaroudi et.al. [23].
@@
2.2.4 Data analysis
The mSSIM, pSNR and CJV serve as quantitative measures for assessing the quality of reconstructions. Additionally, the (unique, in our experiments) tuning parameters $\lambda_{mSSIM},\lambda_{pSNR}$ and $\lambda_{CJV}$ which optimize each metric can be considered. Specifically, $\lambda_{mSSIM}=\arg\max\{mSSIM(\lambda):\lambda\geq 0\}$, $\lambda_{pSNR}=\arg\max\{pSNR(\lambda):\lambda\geq 0\}$ and $\lambda_{CJV}=\arg\min\{CJV(\lambda):\lambda\geq 0\}$. The ratios $\lambda_{mSSIM}/\lambda_{ALM}$, $\lambda_{pSNR}/\lambda_{ALM}$ and $\lambda_{CJV}/\lambda_{ALM}$ express the distance between optimal tuning parameters (with respect to the metrics) and ALMs in terms of the orders of magnitude of the corresponding ratios. We term the ratio $\lambda_{mSSIM}/\lambda$ the magnitude ratio corresponding to $\lambda$ (with respect to mSSIM), and similar notations are reserved for the other metrics. For every pair $({UR}_{\%},{NL}_{\%})$ and every reconstruction, we calculate the mSSIM, pSNR and CJV, alongside their corresponding magnitude rations. Within each fixed pair $({UR}_{\%},{NL}_{\%})$, we employ violin plots to illustrate the distributions of mSSIM, pSNR and CJV across the 50 runs, for analysis. Additionally, shaded error bars for the functions $mSSIM(\lambda/\lambda_{ALM})$, $pSNR(\lambda/\lambda_{ALM})$ and $CJV(\lambda/\lambda_{ALM})$ are incorporated to evaluate the optimality of ALMs as tuning parameters. A value of $\lambda_{mSSIM}/\lambda_{ALM}=1$ signifies that $\lambda_{ALM}$ is the optimal tuning parameter with respect to mSSIM. Similar interpretations apply for the other magnitude ratios. For a comparison, we compute the L-curve tuning parameter $\lambda_{L}$, and execute reconstructions using TV-LASSO with this parameter. Subsequently, we compute $mSSIM(\lambda_{L})$, $pSNR(\lambda_{L})$ and $CJV(\lambda_{L})$, along with the corresponding magnitude ratios pertaining to $\lambda_{L}$ with respect to these three metrics.
2.2.5 The dataset
@@
3 Results
Violin plots for the three metrics across noise levels, undersampling rates and number of runs are displayed in figure 4. The average and standard deviation values of these metrics, as well as their average magnitude ratios, are reported in figure 7.
@@ 3 Results
3.1 Analysis of convergence
Figure 5: Convergence analysis. Subfigure A shows examples of the convergence of ALMA across noise levels and undersampling rates. Each graph corresponds to a fixed couple: noise level and undersampling rate, as illustrated in the legend, showing the iterative progression of the ALM values (vertical axis) at every iteration step (horizontal axis). The dashed lines indicate that the ALM is eventually constant. We observe that the trends displayed by each graph are consistent: the parameter $\lambda^{(k)}$ returned at the first iterations increases, then it descreases until it settles down, with ALMA always stopping in a finite-time. This demonstrates its stability and effectiveness in achieving accurate image reconstructions across noise levels and undersampling rates. Subfigure B shows histograms of the number of iterations needed for ALMA’s convergence across undersampling rates while subfigure C shows the same for different noise levels (see the legends). The number of bins for each histogram is the ceiling of the square root of the number of points, i.e., $\lceil\sqrt{150}\rceil=13$. We observe that the number of iterations for ALMA to stop is mostly concentrated in the interval $[0,10]$ independently of the noise level and the undersampling rate, while the standard deviations increase with the undersampling rate and the noise percentage. This demonstrates that the number of iterations required for ALMA to converge is relatively low even when processing larger amounts of information ($20\%$ sampling) and higher noise percentages ($7\%$ noise).
For fixed noise level and undersampling rate, ALMA performs a reconstruction and updates the ALM approximation at each iteration. We have selected two stopping criteria:
@@ 3.1 Analysis of convergence
Note that the second criterion serves as a convergence criterion for ALMA. This means that if it is satisfied, ALMA not only stops but actually converges (i.e., the sequence $(\lambda^{(k)})_{k}$ of ALMs at step $k$ converges). We observe that in all of our experiments, the sequence of the ALMs is eventually constant. Therefore, in what follows, the term convergence will refer to ALMA stopping due to the fulfillment of the second stopping criterion.
@@
Through empirical analysis, we investigated the convergence of ALMA across various noise levels and undersampling rates. As aforementioned, ALMA converges in a finite time for every undersampling rate and noise level, fig. 5 (A).
@@
The histograms in fig. 5 illustrates the number of iterations required for the convergence of ALMA across different undersampling rates (B) and noise levels (C). Both histograms follow Gaussian distributions that become less concentrated around 0 as the undersampling rate increases (fig. 5 (B)) or as the noise level rises (fig. 5 (C)). In both scenarios, the increase in the number of iterations can be attributed to the amount of information processed by ALMA, which grows with the number of sampling points and the noise level. Let us delve deeper into these cases. On average, ALMA stops after $7.2089\pm 2.9773$ iterations (see fig. 7). For the $10\%$ sampling rate, both the average number of iterations and the standard deviation increase with the noise level. A similar trend is observed for the $15\%$ undersampling rate, as detailed in Table 7. However, the $20\%$ sampling rate exhibits a distinct pattern: the average number of iterations increases from $3\%$ to $5\%$ noise levels and reaches its minimum at a noise level of $7\%$.
3.2 Performance of ALMA with respect to mSSIM
Figure 6: A: Shaded error bar of mSSIM as a function of $\lambda/\lambda_{ALM}$, the shade representing the corresponding standard deviations. The value 1 on the horizontal axis corresponds to $\lambda=\lambda_{\text{ALM}}$.We observe that the corresponding mSSIM values lie on the plateau of the mSSIM graph, just to the right of the maximum. This indicates that the ALM performs nearly optimally with respect to the mSSIM metric. B: Shaded error bar of pSNR as a function of $\lambda/\lambda_{ALM}$. The value 1 on the horizontal axis corresponds to $\lambda=\lambda_{\text{ALM}}$ and the shade represents the corresponding standard deviations. The corresponding points on the graphs are located to the right of the maxima, and this is more pronounced than in the mSSIM case. However, the pSNR of the reconstructions obtained using the ALMs remains high, indicating that the ALMs still produce high-quality reconstructions despite being slightly offset from the optimal point. C: Shaded error bar of CJV as a function of $\lambda/\lambda_{ALM}$. The value 1 on the horizontal axis corresponds to $\lambda=\lambda_{\text{ALM}}$, the shade representing the corresponding standard deviations. The corresponding points on the graphs are located to the right of the maxima, similar to the pSNR case. Nonetheless, the corresponding CJV values are low, indicating that the ALM also performs well with respect to the CJV metric.
@@ 3.2 Performance of ALMA with respect to mSSIM
Images reconstructed utilizing the ALMs computed by ALMA consistently exhibit an average $mSSIM\geq 0.99$, irrespective of the noise level or the undersampling rate (see figure 7). The violin plots for the mSSIM are displayed in figure 4 (color yellow).
@@
Across noise levels and undersampling rates, the tuning parameter that optimizes the mSSIM ($\lambda_{mSSIM}$) is around half of the ALM ($\lambda_{ALM}$), i.e. $\lambda_{mSSIM}\approx 0.52\cdot\lambda_{ALM}$ (see figure 7). However, even if $\lambda_{ALM}$ does not maximize the mSSIM, the points $(1,mSSIM(1))$, which correspond to the mSSIM of the reconstructions obtained with ALMs, always lie on the plateau of the graph of the function $mSSIM(\lambda/\lambda_{ALM})$, indicating that $\lambda_{ALM}$ is in the range of tuning parameters corresponding to almost-optimal mSSIM values (see fig. 6).
3.3 Performance of ALMA with respect to pSNR
@@
Reconstructions obtained using ALMA exhibit high pSNR for every noise level and undersampling rate. As expected the pSNR decreases across noise levels. On average, the pSNR is $\geq 40$dB for reconstructions of images corresponding to $15\%$ and $20\%$ undersampling rates, whereas the average pSNR of reconstructions with $10\%$ undersampling rate is at least 35dB (see figure 7). Other than demonstrating the quality of reconstructions obtained via ALMA in terms of pSNR, this reflects the fact that solving TV-LASSO with ALMs as tuning parameters tend to produce highly regularized images. Again, across noise levels and undersampling rates, the tuning parameter that optimizes the pSNR ($\lambda_{pSNR}$) tends to be approximately 0.47 times the corresponding ALM, i.e. $\lambda_{pSNR}\approx 0.47\cdot\lambda_{ALM}$ (see figure 7). Compared to the graphs $mSSIM(\lambda/\lambda_{ALM})$, the plateau of the functions $pSNR(\lambda/\lambda_{pSNR})$ near their maxima is less pronounced, and the points $(1,pSNR(1))$, corresponding to the pSNR of reconstructions obtained with ALMs, are shifted to the right of their peaks (see fig.6B). In conclusion, ALMs perform almost-optimally with respect to pSNR.
3.4 Performance of ALMA with respect to CJV
@@
Except for the worst-case scenario $UR_{\%}=10\%$, $NL_{\%}=7\%$, the reconstructions obtained by ALMA display CJV values no larger than 0.05, showing that ALMA performs well with respect to CJV as well (see figure 7). The tuning parameter that minimizes the CJV ($\lambda_{CJV}$) is on average 0.45 times the corresponding ALM, i.e. $\lambda_{CJV}\approx 0.45\cdot\lambda_{ALM}$ (see figure 7). However, ALMs still perform almost optimally with respect to CJV, and the points $(1,CJV(1))$, corresponding to the CJV of reconstructions obtained with ALMs, are shifted to the right of their minima, in accordance with the behavior of the graphs $pSNR(\lambda/\lambda_{CJV})$ (see fig.6C).
3.5 Comparison with the L-curve parameter